OpenStudy (anonymous):

Calculate the equation of the line that passes throught the points (1,5) and (3,6)

OpenStudy (accessdenied):

Do you know the general equation for lines?

OpenStudy (anonymous):

we have been using the equations y+a(x-h)+k or y=(x-h)+k

OpenStudy (anonymous):

I answered a similar question a while ago: http://openstudy.com/study#/updates/50cbc57de4b0031882dbd6c4 Now that I think about it I answer this every once in a while on Open Study. Maybe the teachers should do a better job explaining... o_0

OpenStudy (anonymous):

thank you i think i know what im doing now(:

OpenStudy (accessdenied):

If we want to use that equation, I would insert one point in for (h, k) and solve for a by substituting in the other as (x, y). y = a(x - h) + k We can let (h, k) be either point and (x, y) be the other point, it seems. We then just solve for the value of a to use in the equation.

OpenStudy (anonymous):

okay thank you (:

OpenStudy (accessdenied):

You're welcome! :)

OpenStudy (anonymous):

i have one other question, how do you calculate the slope and distance between a line?

OpenStudy (accessdenied):

Hmm... could you explain the "distance of a line"? As in, the distance between a line and a point? Or between two points on the line? Slope is just calculated as the ratio of the differences in y values and x values of two points on a line. Or in other words, how much it 'rises' over how much it 'runs. m = (y2 - y1)/(x2 - x1) from two points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

the distance between two points like from point to point

OpenStudy (accessdenied):

The general formula of the distance between two points is just \( \displaystyle d = \sqrt{ (x_2 - x_1)^2 + (y_2 - y_1)^2} \) It comes from Pythagorean theorem, where the graph is like this: